OpenStudy (anonymous):

write an equation for the line that passes through (-2,7) and (3,-8)

Mehek (mehek14):

ok so find the slope \(\dfrac{-8 - 7}{3-(-2)}=\)

OpenStudy (anonymous):

thank you so much

Mehek (mehek14):

that's not the whole thing

OpenStudy (anonymous):

how i do that?

OpenStudy (anonymous):

finding the slope

Mehek (mehek14):

first find the slope then find y-intercept

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so its -15/5?

OpenStudy (anonymous):

hello???

OpenStudy (sdfgsdfgs):

y=mx+a is called the slop-intercept form of a line; where m is the slope and a is the y-intercept. but i will probably do it in a slightly different way. u got the slope already; would u plz simplify it to a single number instead of a fraction?

OpenStudy (anonymous):

ok can you plz help me im a senior and i have to flash back to freshmen year

OpenStudy (anonymous):

i forgot this stuff

OpenStudy (sdfgsdfgs):

hahahaa okok being a senior is hard enough n freshman yr seems like ancient time right ;) OK so what the slope simplifies to plz?

OpenStudy (anonymous):

yes!!!!!!!!

OpenStudy (anonymous):

you understand

OpenStudy (sdfgsdfgs):

so the slope is....

OpenStudy (anonymous):

3

OpenStudy (sdfgsdfgs):

-15/5?

OpenStudy (sdfgsdfgs):

missing something there?

OpenStudy (anonymous):

oh yeah -3

OpenStudy (sdfgsdfgs):

OK let (x,y) be the variable of ur eqn...since (3,-8) is a point on that eqn, slope=(y-(-8))/(x-3) since slope=-3, (y-(-8))/(x-3)=-3 can u simplify n re-arrange the eqn above? dats ur ans :)

OpenStudy (anonymous):

thx!

OpenStudy (sdfgsdfgs):

heheheee thanx 4 da medal! if u want, plz post ur ans, i will be happy to double-check 4 u :)